The atomic density of a solid is 5.85 ×1028 m−3. Its electrical resistivity is 1.6 ×10−8Ωm. Assume that each atom contributes one conduction electron:
The relaxation time (mean free time), in seconds, of the conduction electrons is?
1
Expert's answer
2021-09-19T09:25:23-0400
"\tau=\frac{\sigma m}{ne^2}\\\tau=\frac{(9.11\cdot10^{-31})}{(1.6\cdot10^{-8})(5.85\cdot10^{28})(1.6\cdot10^{-19})^2}=3.8\cdot10^{-14}\ s"